OpenStudy (anonymous):

Find two sets of polar coordinated for (1,1) for zero is less than or equal to theta is less than two pie

OpenStudy (anonymous):



OpenStudy (anonymous):

you need \(r\) and \(\theta\) \(r=\sqrt{1^2+1^2}=\sqrt2\) by pythagoras, and \(\theta\) should be more or less obvious